# Fermi-LAT Discovery of a Gamma-ray Outburst from the Peculiar Compact Steep Spectrum Radiogalaxy 3C 216

### v1

@@ Abstract
3C 216 is an extra-galactic radio source classified as a compact steep spectrum (CSS) object, associated with the source 4FGL J0910.0+4257 detected by the Large Area Telescope (LAT) on board the Fermi Gamma-ray Space Telescope. The source exhibits extended radio structures as well as an inner relativistic jet. In general, jets accelerated by active galactic nuclei (AGNs) are efficient sources of non-thermal radiation, spanning from the radio band to X-ray and gamma-ray energies. Due to relativistic beaming, much of this radiation, particularly in the high-energy domain, is concentrated within a narrow cone aligned with the jet’s direction. Consequently, high-energy emission is more easily detected in blazars, where the jet is closely aligned with the line of sight of the observer. Beginning in November 2022 , Fermi-LAT observed increased gamma-ray activity from 3C 216, culminating in a strong outburst in May 2023. This event was followed up by observations from the Neil Gehrels Swift Observatory telescope. In this work, we perform a careful analysis of the multifrequency data (gamma-ray, X-ray, UV, optical) collected during this observational campaign. We find that the spectral energy distribution of the flaring source evolves in a coherent way, suggesting a common origin for the multifrequency emission. These results support the interpretation of the gamma-ray emission within a single zone synchrotron self-Compton (SSC) model, with important implications for the mechanisms powering high-energy radiation in AGN jets.
@@ 1 Introduction
The active galactic nucleus (AGN) 3C 216 ($z=0.670$, (Smith & Spinrad, 1980)) is classified as an extragalactic CSS radio source. On arcsecond scales, it consists of a central component surrounded by a more extended structure with an angular size of $4.5^{\prime\prime}$, corresponding to a projected linear size (LS) of $56$ kpc (Principe et al., 2021). Although its radio spectrum peaks at low frequency ($\nu<0.5$ GHz), an upturn is observed at a few GHz, attributed to the presence of a central flat-spectrum core (Taylor et al., 1995). This core exhibits a significant misalignment with the outer structure, along with superluminal jet component motion, with a velocity of approximately $\sim 4c$ (Venturi et al., 1993; Paragi et al., 2000). While the compact nature of CSS sources is typically attributed to their young age within an evolutionary framework (Fanti et al., 1995), the characteristics of the central component, a pronounced optical polarization and variability (Impey et al., 1991), strongly suggest the presence of a blazar core.
The Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope, is a pair-conversion telescope designed to detect photons in the energy range from $20$ MeV to $2$ TeV. The LAT observed a pronounced enhancement in the gamma-ray activity from the direction of the source on 2022-12-08 (La Mura, 2022) and on 2023-05-01 (Giacchino et al., 2023). Multifrequency observations of the source have been carried out in order to associate the gamma-ray flare with activity at different wavelengths. The observations across various bands provide insights into both the origin of the gamma-ray emission and the structure of the radio jet. As we will demonstrate in this work, during the flare emission, the source exhibited a hard gamma-ray spectrum, in contrast to its typically soft spectrum during quiescent periods. This behavior can be explained by a synchrotron self-Compton (SSC) model across all frequencies, with evidence of a cooling process in the days after the flare peak.
In the Sec. 2 we report the main results of the Fermi-LAT observations of the flaring activity that occurred in May 2023 and in the Subsec. 2.1 we provide a refined analysis of the Fermi-LAT data, highlighting the main features of the source in terms of its gamma-ray spectrum and light curve. We then describe in Sec. 3 the Swift ToO observations of 3C 216, with a detailed analysis presented in Subsec. 3.1. Finally, the discussion and conclusion are provided in Sec. 4.
@@ 2 Fermi-LAT observations
The CSS source 3C 216 is located at R.A. = 137.38957 deg, DEC. = 42.89624 deg [J2000,  (Petrov et al., 2005)]. It is associated with the 4FGL J0910.0+4257 source in the third data release of 4FGL catalog (4FGL-DR3, Abdollahi et al., 2022) with coordinates R.A. = $137.51\pm 0.11$ deg, DEC. = $42.96\pm 0.11$ deg.
Fermi-LAT detected flaring activity from the direction of this source. A preliminary Fermi-LAT analysis of the Flare Advocate data for 2023-05-01 revealed that the daily-averaged gamma-ray flux increased by a factor of $177$ relative to the average flux reported in the 4FGL-DR3, $\langle\Phi\rangle_{\gamma}=(7.5\pm 1.4)\times 10^{-9}\,\mathrm{ph\,cm^{-2}\,s^{-1}}$ in the $1-100$ GeV energy range. The photon index decreased from the 4FGL averaged value, ($2.57\pm 0.10$), to ($2.11\pm 0.09$). This represented both the highest daily flux and the hardest photon index recorded for this source by Fermi-LAT to date (Giacchino et al., 2023). Due to the exceptional magnitude of this flare, the source association has been confirmed through a dedicated follow-up analysis with Fermi-LAT, complemented by multifrequency observations with Swift, which we will describe in detail later.
@@ 2.1 Gamma-ray analysis
The data analysis was performed with Fermipy [^1] https://fermipy.readthedocs.io/en/latest/ (version 1.2.0.) (Wood et al., 2017) and the Fermitools [^2] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ software packages (version 2.2.0). We analyzed three periods: all Fermi-LAT data up to one month after the event (see Table 1 for details), the flare period, defined as the epoch during which Fermi-LAT detected an increasing activity trend on a weekly time scale, and the peak flare (PF) period, covering the maximum daily activity and the subsequent multi-frequency follow-up granted by the requested Swift observations.
@@
The time range chosen for flare period is $2022$-$11$-$14$ to $2023$-$06$-$06$, while the PF period goes from $2023$-$04$-$28$ to $2023$-$05$-$11$. The energy range selected for the analysis is $[0.1-300]$ GeV, using Pass 8 events (Atwood et al., 2013) and all the available photons of the SOURCE class, excluding those arriving with zenith angles greater than $90^{\circ}$ for energies smaller than $1$ GeV otherwise greater than $105^{\circ}$. Moreover, we choose a region of interest (ROI) of $15^{\circ}\times 15^{\circ}$ around our target, with a pixel size of $0.1^{\circ}$ and $8$ evenly spaced logarithmic energy bins. We utilize the P8R3_SOURCE_V3 instrumental response functions (IRFs), along with the galactic diffuse model gll_iem_v07.fits and the isotropic diffuse model P8R3_SOURCE_V3_v1.txt. The model used for the analysis includes all sources in the 4FGL-DR3 catalog [^3] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html located at a distance $\leq 20^{\circ}$ from 4FGL J0910.0+4257. A summary of the parameters is reported in Table 1. In the spectral analysis we left the model parameters of the isotropic and diffuse background as well as the sources within $3^{\circ}$ of our target free to vary. To produce the spectral energy distribution (SED) of all the three periods, we used eight logarithmically spaced bins between $100$ MeV and $10$ GeV and four selected energy bins from $10$ GeV and $300$ GeV obtaining the plot in Fig. 1. The plot clearly shows that the PF (blue) and flare (green) period fluxes are much higher than the values obtained considering all the Fermi-LAT (red) observations, confirming an interesting enhanced activity.
@@
Figure 1: SED of 3C 216 for three periods: in blue PF (Peak Flare) period, in green flare period, in red total Fermi-LAT period. The upper limit is reported when $TS\leq 10$
Period $\Gamma$ $\Phi_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$ $\alpha$ $\beta$ $N_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$
PF $1.97\pm 0.05$ $(5.87\pm 0.46)\times 10^{-11}$ $1.79\pm 0.23$ $0.12\pm 0.09$ $(3.83\pm 0.72)\times 10^{-11}$
Flare $2.03\pm 0.01$ $(2.45\pm 0.09)\times 10^{-11}$ $1.99\pm 0.16$ $0.06\pm 0.03$ $(1.73\pm 0.25)\times 10^{-11}$
Total $2.17\pm 0.04$ $(1.69\pm 0.07)\times 10^{-12}$ $1.95\pm 0.20$ $0.12\pm 0.07$ $(1.02\pm 0.47)\times 10^{-12}$
DR4 $2.43\pm 0.07$ $(8.05\pm 0.71)\times 10^{-13}$ $2.24\pm 0.15$ $0.17\pm 0.10$ $(9.02\pm 0.91)\times 10^{-13}$
@@
Table 2: List of averaged spectral parameters for three periods (PF, flare, total): photon index for power-law $\Gamma$, normalization factor for power-law $\Phi_{0}$ ($\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}$), slope of the log-parabola $\alpha$, curvature log-parabola $\beta$, normalization factor for log-parabola $N_{0}$ ($\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}$). The 4FGL-DR4 values are reported in the catalog (Ballet et al., 2023).
@@
AGN gamma-ray spectra are generally well represented by either a log-parabola (LP) function or by a power-law (PL) one. The LP function is described as:
@@
where the normalization factor $\Phi_{0}$ (MeV-1 cm-2 sec-1) is the flux density at $E_{0}$, $\alpha$ is the slope index, $\beta$ is the curvature, and $E_{0}$ (MeV) is the scale parameter.
The PL function, instead, is described as:
@@
where the spectral parameters are the normalization factor $N_{0}$ in (MeV-1 cm-2 sec-1), spectral index $\Gamma$ and the energy scale $E_{b}$ (MeV).
The difference in the spectral state obtained from the analysis of the three considered periods can be appreciated by the comparison of their PL indices and LP spectral parameters with the corresponding values reported in the 4FGL-DR4 catalog  (Ballet et al., 2023), as shown in Table 2. The flare period is characterized by a spectral hardening that is detected as a decrease of power-law index and log-parabola slope, which becomes more striking at the time of peak flux. According to  Abdollahi et al. (2022), the sources are represented with a curved spectral model when $TS_{curv}=2[log\mathcal{L}_{curv}-log\mathcal{L}_{PL}]>4\,(2\sigma)$. In our case the curved spectral model is a LP, and all the three periods are characterized by a larger $TS_{curv}$ than threshold: $TS_{LP}=3.0\sigma$ for the total period, $TS_{LP}=3.2\sigma$ for the flare period, and $TS_{LP}=3.1\sigma$ for the peak flare period. Notably, the 4FGL-DR3 catalog reports a curvature significance of $TS_{LP}\sim 1.68\sigma$ for this source and the preferred spectral model is a power-law. On the contrary, the longer monitoring of 4FGL-DR4 (Ballet et al., 2023) points to a curvature significance of $TS_{LP}\sim 2.39\sigma$, supporting a curved spectrum, in agreement with our result. Likely the flaring activity increased the number of photons collected from this source allowing a more accurate reconstruction of its spectral form. For this reason, the spectrum during the flaring period is well represented by a log-parabola model, as shown in Fig. 2.
@@
Figure 3: Fermi-LAT light curves of 3C 216 for the three periods considered here with different binnings: in the top panel (blue) Peak Flare period daily binned, with dashed lightblue lines representing the days of Swift detection (2023-05-03, 04, 06, 08, 09); in the middle panel (green) flare period weekly binned, with the dotted orange line representing the first flare ($2022$-$12$-$08$) and the dashed orange line the beginning of the PF activity ($2023$-$05$-$01$), the blue area refers to the time range of the top panel; in the bottom panel (red) the total period with six months binning, the green and blue areas refer to the middle and top pannel, respectively. Upper limits are reported when $TS\leq 10$.
We also extracted the light curves, by allowing the flux normalization of the sources within $3^{\circ}$ of the target to vary freely while freezing all the other parameters from the baseline analysis, in Fig 3. The top panel displays the flux around the PF period using $1$-day time bins from $2023$-$04$-$28$ to $2023$-$05$-$10$. The middle panel shows the flare period with $1$-week time bins, and the bottom panel shows the total period with $6$-month time bins. The analysis of the flare period reveals that the flare activity achieved a first maximum on 2022-12-8, marked by the orange dotted line in the middle panel, followed by an epoch of moderate activity before the exceptional outburst occurring on 2023-05-01, highlighted by the orange dashed line. The light curve in PF the period illustrates the sharp rise of the flux leading to the flare peak (top panel of Fig. 3). The blue dashed lines indicate the Swift observations conducted in 2023.
@@
We studied the spectral variability in relation to the flux during the flaring activity, as shown in the top panel of Fig. 4. In this short time range, the spectral curvature is not particularly significant, and we can use a PL model to obtain a more accurate evaluation of the spectral variability. This analysis aims to characterize the synchrotron emission model during the flare and its subsequent cooling phase. The plot shows the evolution of the flare’s spectral parameters from $2023$-$04$-$30$ at $18$:$00$:$00$ to $2023$-$05$-$04$ at $18$:$00$:$00$, in $12$-hour bins, following the flux increase that led to the maximum observed gamma-ray daily emission and the duration of the subsequent Swift multi-frequency follow-up. The observations are numbered to represent their chronological order. The chronological evolution is suggestive of an anticlockwise trail for the first four bins (numbered from 0 to 3), while the spectral index appears to become more stable at later times. This is consistent with the expectations implied by a single-zone SSC scenario, where the radiative efficiency and the energy distribution of the emitting particles are subject to the competing effects of a fast energy injection, occurring nearly simultaneously for all particles, and a radiative cooling that affects high energy particles more quickly than low energy ones  (Kirk et al., 1998). In the bottom panel of Fig. 4, we associated the light curve with the corresponding spectral data points to illustrate the flux variability over time.
@@
Figure 4: Top panel: the gamma-ray photon index of 3C 216 as a function of the gamma-ray flux. Bottom panel: the light curve of 3C 216. In both panels, the numbers indicate the chronological order of emissions, with the starting point at $2023$-$04$-$30$ at $18$:$05$:$49$ designed as number 00, and the finishing point at $2023$-$05$-$04$ at $18$:$05$:$49$ designed as number $8$. The spectral index is evaluated in a 121212\,12hr binned light curve.

### v2

@@ Abstract
3C 216 is an extra-galactic radio source classified as a compact steep spectrum (CSS) object, associated with the source 4FGL J0910.0+4257 detected by the Large Area Telescope (LAT) on board the Fermi Gamma-ray Space Telescope. The source exhibits extended radio structures as well as an inner relativistic jet. In general, jets accelerated by active galactic nuclei (AGNs) are efficient sources of non-thermal radiation, spanning from the radio band to X-ray and gamma-ray energies. Due to relativistic beaming, much of this radiation, particularly in the high-energy domain, is concentrated within a narrow cone aligned with the jet’s direction. Consequently, high-energy emission is more easily detected in blazars, where the jet is closely aligned with the line of sight of the observer. Beginning in November 2022, Fermi-LAT observed increased gamma-ray activity from 3C 216, culminating in a strong outburst in May 2023. This event was followed up by observations from the Neil Gehrels Swift Observatory telescope. In this work, we perform a careful analysis of the multifrequency data (gamma-ray, X-ray, UV, optical) collected during this observational campaign. We find that the spectral energy distribution of the flaring source evolves in a coherent way, supporting a common origin for the multifrequency emission. These results suggest that the spectral energy distribution (SED) observed during the outburst was dominated by a single emission zone, where synchrotron self-Compton processes (SSC) played a primary role. Since single-zone SSC models have typically less free parameters than multi-zone alternatives, they are a powerful probe of the physical conditions of the high-energy emitting regions. Therefore, observing SSC radiation even in CSS sources improves our understanding of the production of high-energy radiation in AGN jets.
@@ 1 Introduction
The active galactic nucleus (AGN) 3C 216 ($z=0.670$, Smith & Spinrad 1980) is classified as an extragalactic CSS radio source. On arcsecond scales, it consists of a central component surrounded by a more extended halo structure with an angular size of $4.5^{\prime\prime}$, corresponding to a projected linear size (LS) of $56$ kpc (Principe et al., 2021). Although its radio spectrum peaks at low frequency ($\nu<0.5$ GHz), an upturn is observed at a few GHz, attributed to the presence of a central flat-spectrum core plus a faint halo component (Taylor et al., 1995). The core exhibits a significant misalignment with the outer structure, along with superluminal jet component motion, with a velocity of approximately $\sim 4c$ (Venturi et al., 1993; Paragi et al., 2000). While the compact nature of CSS sources is typically attributed to their young age within an evolutionary framework (Fanti et al., 1995), the characteristics of the central component, a pronounced optical polarization and its variability (Impey et al., 1991) strongly suggest the presence of a blazar core.
The Large Area Telescope (LAT, Atwood et al., 2009) onboard the Fermi Gamma-ray Space Telescope, is a pair-conversion telescope designed to detect photons in the energy range from $20$ MeV to $2$ TeV. The LAT observed a pronounced enhancement in the gamma-ray activity from the direction of the source on 2022-12-08 (La Mura, 2022) and on 2023-05-01 (Giacchino et al., 2023). Multifrequency observations of the source were carried out, thanks to a Swift Target of Opportunity (ToO) request, in order to associate the $\gamma$-ray flare with activity at different wavelengths. The observations across various bands provide insights into both the origin of the $\gamma$-ray emission and the structure of the relativistic jet. During the flare, the source exhibited a hard gamma-ray spectrum, as opposed to the typically soft one observed in quiescent periods, together with coherent variability from low energies up to $\gamma$ rays. This behavior can be explained in terms of a single-zone emission model, where it is likely that synchrotron self-Compton radiation dominated the emission across all frequencies during the outburst, with evidence of a cooling process in the following days.
The paper is structured as follows: in Sec. 2 we present the detection of the exceptional gamma-ray activity observed by Fermi-LAT in May 2023, with a detailed analysis of the Fermi-LAT data and its comparison with long term observations given in Sec. 2.1; in Sec. 3 we describe the Swift ToO observations of 3C 216, detailing their analysis in Sec. 3.1; in Sec. 4 we discuss our results and, finally, in Sec. 5 we summarize our conclusions.
@@ 2 Fermi-LAT observations
3C 216 is located at $R.A.=137.38957\,$deg, $DEC.=42.89624\,$deg (J2000, Petrov et al., 2005). It is associated with the $\gamma$-ray source 4FGL J0910.0+4257, with coordinates $R.A.=(137.51\pm 0.11)\,$deg, $DEC.=(42.96\pm 0.11)\,$deg in the third release of the 4FGL catalog (4FGL-DR3, Abdollahi et al., 2022). In the course of its regular monitoring of the sky, Fermi-LAT detected an increase of activity from the direction of this source starting from the mid of November 2022. On 2023-05-01, a quick-look analysis showed that the source exhibited a sudden outburst, achieving a daily averaged $\gamma$-ray flux of $\langle\Phi\rangle_{daily}=(1.32\pm 0.15)\times 10^{-6}\,\mathrm{ph\,cm^{-2}\,s^{-1}}$, a factor 176 times higher than the average flux $\langle\Phi\rangle_{\gamma}=(7.5\pm 1.4)\times 10^{-9}\,\mathrm{ph\,cm^{-2}\,s^{-1}}$ reported in the 4FGL-DR3 catalog over the $0.1-100\,$GeV energy range. The photon index decreased from the catalog value of $(2.52\pm 0.10)$ to $(2.11\pm 0.09)$. This event represented both the highest daily flux and the hardest spectral state reported for the source to date (Giacchino et al., 2023).
Due to the exceptional magnitude of the event, multi-frequency follow-up observations were requested by activating a Swift ToO, which detected 3C 216 in a flaring state in optical, UV, and X-rays, confirming its association with 4FGL J0910.0+4257. We therefore analyzed the $\gamma$-ray history of the source and its associated X-ray and optical/UV observations to characterize the origin of its flaring activity.
[…]
@@ 2.1 Gamma-ray analysis
In order to characterize the status of 3C 216 during the outburst and to compare it with its long-term averaged properties, we analyzed the data collected by Fermi-LAT over three distinct periods. The first, which we refer to as the total monitoring period, covers all the Fermi-LAT observations of the source, from the start of science operations on 2008-08-04 up to 2023-06-06, approximately one month after the main $\gamma$-ray outburst. The second, which we call the flare period, corresponds to the start of the weekly rising trend in the activity of the source, on 2022-11-14, up to the end of the total period. The third one, which we call the flare peak, extends from 2023-04-28 to 2023-05-10, covering the time range corresponding to the highest level of daily $\gamma$-ray activity and the subsequent Swift follow-up campaign. For all these periods we performed a light curve analysis, plotted in Fig. 1, and a standard binned analysis.
Table 1: Table of fermipy analysis parameters.
@@
In our study, we used fermipy v1.2.0111https://fermipy.readthedocs.io/en/latest/ (Wood et al., 2017) and the FermiTools v2.2.0.222https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ We selected all the observations in the $[0.1-300]\,$GeV energy range, using Pass 8 events (Atwood et al., 2013) and all the available photons of the SOURCE class, excluding those arriving with zenith angles greater than $90^{\circ}$ for energies smaller than $1$ GeV, otherwise greater than $105^{\circ}$. Moreover, we chose a region of interest (ROI) with an aperture radius of $15^{\circ}$ around the target, with a pixel size of $0.1^{\circ}$ and $8$ evenly spaced logarithmic energy bins. We utilized the P8R3_SOURCE_V3 instrumental response functions (IRFs), along with the galactic diffuse model gll_iem_v07.fits and the isotropic diffuse model P8R3_SOURCE_V3_v1.txt. The model used for the analysis includes all sources in the 4FGL-DR3 catalog [^3] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html located at a distance $\leq 20^{\circ}$ from 4FGL J0910.0+4257. A summary of the analysis parameters is reported in Table 1.
@@
Figure 2: SED of 3C 216 for three periods: in flare peak (blue), in flare (green), and total (red) periods. Upper limits are reported when $TS\leq 10$.
For the spectral analysis we left the model parameters of the isotropic and diffuse background as well as the sources within $3^{\circ}$ of our target free to vary. To produce the spectral energy distribution (SED) of all the three periods, we used eight logarithmically spaced bins between $100$ MeV and $10$ GeV and four energy bins from $10$ GeV and $300$ GeV, obtaining the plot in Fig. 2. The result clearly shows that the flare peak (blue) and the flare (green) period fluxes are much higher than the values obtained considering all the Fermi-LAT observations (red), confirming an interesting enhanced activity. For the light curve analysis shown in Fig. 1, instead, we allowed the flux normalization of the sources within $3^{\circ}$ of the target to vary freely while freezing all the other parameters from the baseline analysis. The top panel displays the total period with $6$-month time bins, the middle panel shows the flare period in a weekly binning, and the botton panel presents the evolution leading to the flare peak and the subsequent follow-up, using $1$-day time bins from $2023$-$04$-$28$ to $2023$-$05$-$10$. For the total period we have identified statistically significant variations compared to the 4FGL-DR3 averaged flux with the Bayesian Blocks algorithm (Scargle et al., 2013) as implemented in astropy [^4] https://docs.astropy.org/en/stable/api/astropy.stats.bayesian_blocks.html.
The analysis of the flare period reveals that the flare activity achieved a first maximum on 2022-12-8, marked by the orange dotted line in the middle panel of Fig. 1,followed by an epoch of moderate activity before the exceptional outburst occurring on 2023-05-01, highlighted by the orange dashed line. The blue curve in the peak period illustrates the sharp rise of the flux leading to the flare peak (bottom panel of Fig. 1), while the light blue dashed lines indicate the Swift observations conducted during the follow-up campaign. The maximum flux obtained by the analysis is $\Phi_{max}=(1.22\pm 0.13)\times 10^{-6}$ ph cm-2 s-1 with a photon index of $2.06\pm 0.12$. The shaded areas in the top and middle panels refer to the time range of the following ones.
@@
Table 2: List of averaged spectral parameters for three periods (flare peak, flare, total): photon index for power-law $\Gamma$, normalization factor for power-law $\Phi_{0}$, slope of the log-parabola $\alpha$, curvature log-parabola $\beta$, normalization factor for log-parabola $N_{0}$. We include, for reference, also the values reported in the 4FGL-DR4 catalog (Ballet et al., 2023)
. Period $\Gamma$ $\Phi_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$ $\alpha$ $\beta$ $N_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$ Flare Peak $1.97\pm 0.05$ $(5.87\pm 0.46)\times 10^{-11}$ $1.79\pm 0.23$ $0.12\pm 0.09$ $(3.83\pm 0.72)\times 10^{-11}$ Flare $2.03\pm 0.01$ $(2.45\pm 0.09)\times 10^{-11}$ $1.99\pm 0.16$ $0.06\pm 0.03$ $(1.73\pm 0.25)\times 10^{-11}$ Total $2.17\pm 0.04$ $(1.69\pm 0.07)\times 10^{-12}$ $1.95\pm 0.20$ $0.12\pm 0.07$ $(1.02\pm 0.47)\times 10^{-12}$ 4FGL-DR4 $2.43\pm 0.07$ $(8.05\pm 0.71)\times 10^{-13}$ $2.24\pm 0.15$ $0.17\pm 0.10$ $(9.02\pm 0.91)\times 10^{-13}$
2.2 $\gamma$-ray spectral properties
@@
AGN gamma-ray spectra are generally well represented by either a log-parabola (LP) function or by a power-law (PL) one. The LP function is defined as:
@@
where the normalization factor $\Phi_{0}$ (MeV-1 cm-2 sec-1) is the flux density at $E_{0}$, $\alpha$ is the spectral index, $\beta$ is the curvature, and $E_{0}$ (MeV) is the scale parameter. The PL function, instead, is defined as:
@@
where the parameters are the normalization factor $N_{0}$ in ($\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}$), the spectral index $\Gamma$ and the energy scale $E_{b}$ (MeV). According to  Abdollahi et al. (2022), the sources are represented with a curved spectral model when $TS_{curv}=2[log\mathcal{L}_{curv}-log\mathcal{L}_{PL}]>4\,(=2\sigma)$. In our case the likelihood for curvature is evaluated by testing LP against PL spectral models. All the three periods are characterized by a larger $TS_{curv}$ than threshold: $TS_{LP}=3.0\sigma$ for the total period, $TS_{LP}=3.2\sigma$ for the flare period, and $TS_{LP}=3.1\sigma$ for the flare peak period. Notably, the 4FGL-DR3 catalog reports a curvature significance of $TS_{LP}\sim 1.68\sigma$ for this source and the preferred spectral model is a power-law. On the contrary, the longer monitoring of 4FGL-DR4 (Ballet et al., 2023) points to a curvature significance of $TS_{LP}\sim 2.39\sigma$, supporting a curved spectrum, in agreement with our result. Likely the flaring activity increased the number of photons collected from this source allowing a more accurate reconstruction of its spectral form. As a consequence, the spectrum during the flaring period is well represented by a log-parabola model, as shown in Fig. 3.
@@
The evolution of the spectral state observed for the total period, the flare period and the flare peak period can be appreciated by looking at the corresponding spectral index and curvature parameters, listed in Table 2, in comparison with the values reported in 4FGL-DR4 (Ballet et al., 2023). We see from these results that the flare period is characterized by a spectral hardening, with respect to the average state of the source, that is detected as a decrease of the power-law index and of the log-parabola slope, becoming more evident at the time of the peak flux.
@@
To further characterize the spectral variability during the flare peak, we studied the spectral variability in relation to the flux, performing a 121212\,12hr binned analysis from 2023-04-30 to 2023-05-03 (Fig. 4). In this short time range, the spectral curvature is not particularly significant, and we can use a PL model to obtain a more accurate evaluation of the spectral variability. This analysis aims to characterize the spectral behaviour during the main $\gamma$-ray outburst and the subsequent decreasing stage. The observations are numbered to represent their chronological order. In spite of the rather large uncertainties, the evolution shown in the upper panel is suggestive of an anticlockwise trail for the first four bins (numbered from 0 to 3), while the spectral index appears to become more stable at later times. This is consistent with the expectations implied by a single-zone SSC scenario, where the radiative efficiency and the energy distribution of the emitting particles are subject to the competing effects of a fast energy injection, occurring nearly simultaneously for all particles, and a radiative cooling that affects high energy particles more quickly than low energy ones  (Kirk et al., 1998). In the bottom panel of Fig. 4, we associated the light curve with the corresponding spectral data points to illustrate the flux variability over time.
@@
Figure 4: Top panel: the gamma-ray photon index of 3C 216 as a function of the gamma-ray flux. Bottom panel: the light curve of 3C 216. In both panels, the numbers indicate the chronological order of the observations, extracted as 121212\,12hr bins starting from $2023$-$04$-$30$ at $18$:$05$:$49$ (designed as number 00) and finishing point on $2023$-$05$-$04$ at $18$:$05$:$49$ (designed as number $8$).